OpenStudy (anonymous):

slope=4 passing through (1,3) use the given conditions to write an equation in the point slope formand and the slope intercept form

OpenStudy (anonymous):

y-3=4(x-1)

OpenStudy (anonymous):

can you explain

OpenStudy (anonymous):

y=4x-1

OpenStudy (anonymous):

point slope equation is : y-y1=m(x-x1), where m is the slope and (x1,y1) is a given point in your case (1,3). The slope intercept form is simply the point slope form simplified

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

can you walk me through the next one

OpenStudy (anonymous):

slope=8 passing through (4,-1)

OpenStudy (anonymous):

y-y1=m(x-x1) y-(-1)=8(x-4) y+1=8(x-4): point slope form

OpenStudy (anonymous):

y+1=8(x-4) y+1=8x-32 y=8x-33

OpenStudy (anonymous):

that was slope intercept form